OpenStudy (anonymous):

Solve the following equation. Please show all of your work. 7+5/7 x=-13

OpenStudy (anonymous):

(7+5/7x) -7 = -13 -7 5/7x = -20 7* (5/7x) = -20 * 7 (5x) /5 = -140 /5 x= -24

OpenStudy (anonymous):

thanks!